OpenStudy (anonymous):

when solving for y=-(2/3)x+30....you did you get -30*3? need help in the steps. PLEASE!

OpenStudy (anonymous):

trying to solve for intercepts x and y I don't upstand steps.

myininaya (myininaya):

To find x-intercept set y=0 and solve for x; y=-(2/3)x+30 0=-(2/3)x+30 -30 -30 -30=-(2/3)x *(-3/2) *(-3/2) (-30)(-3/2)=x x=(-30)(-3/2) x=(-30/2)(-3) x=(-15)(-3) x=45

myininaya (myininaya):

To find y-intercept set x=0 and solve for y; y=-(2/3)x+30 y=-(2/3)(0)+30 y=0+30 y=30

myininaya (myininaya):

so x-intercept is (45,0) y-intercept is (0,30)

OpenStudy (anonymous):

thanks